OpenStudy (gabylovesyou):

If A = {positive integers divisible by 2} and B = {integers}, what is A ∩ B? { } {2, 4, 6, 8 …} {2, 4, 6, 8, 10} {1, 2, 3, 4, 5 …}

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

@TheMind No. @Gabylovesyou do you know what the intersection operator means?

OpenStudy (gabylovesyou):

no i don't @vf321

OpenStudy (shaik0124):

{2, 4, 6, 8 …}

OpenStudy (anonymous):

Okay. The intersection operator \(\cap\) means, quite literally, the intersection of the venn diagrams which represent each set that's being operated on.

OpenStudy (shaik0124):

intersection indicates comon elements among the two sets so

OpenStudy (anonymous):

So \(\left\{ 1,2,3 \right\}\cap\left\{ 3,4,5\right\}=\left\{ {3} \right\}\)

OpenStudy (anonymous):

Do you see why? If you draw the Venn diagrams,  3 is the only one in the middle.

OpenStudy (shaik0124):

a={2,4,6,8,10,12,14....} b={1,2,3,4,5,6,7...} a intersection b={2,4,6,8,10....}

OpenStudy (anonymous):

@shaik0124 how so?

OpenStudy (shaik0124):

intersection gives common elments among two sets

OpenStudy (shaik0124):

kk @vf321

OpenStudy (anonymous):

@Gabylovesyou if your question was answered please close the question (top-left corner, under "Mathematics")

OpenStudy (gabylovesyou):

it wasnt @vf321 i still dont know which one

OpenStudy (anonymous):

OKay lets work through this. Do you have an understanding of what an intersection is now?

OpenStudy (gabylovesyou):

yes

OpenStudy (anonymous):

Okay. Do you agree that the set of all integers looks like this:\[\left\{...-2,-1,0,1,2,...\right\}\]?

OpenStudy (gabylovesyou):

yessssss

OpenStudy (anonymous):

okay. Good. Now I have a quick question for you. Do you agree that "all positive integers divisible by 2" are also integers? 'Cause that means that the set which contains the former is a SUBSET of the latter.

OpenStudy (anonymous):

Or, in Venn diagrams, with A = All integers, and B = All positive ints divisible by 2, do you see why: